A prism whose refracting angle is 600 causes a minimum deviation of 48 in a monochromatic bean of light. Calculate the refractive index of the prism for that wavelength.


1
Expert's answer
2021-07-02T09:39:01-0400

"\mu=\frac{\sin(\frac{A+\delta}2)}{\sin \frac A2},"

"\mu=\frac{\sin(\frac{60°+48°}2)}{\sin \frac {60°}2}=\frac{\sqrt 5+1}2=1.618."